OpenStudy (katielong):

[A-level C3 Calculus] differentiate y=(6x-3)(2x-1)^5 Please Help!

OpenStudy (kainui):

You can take the natural log of both sides, and take the derivative that way.

OpenStudy (katielong):

okay how would i do that? the text book only mentions the quotient, chain and product rules though..?

OpenStudy (nincompoop):

can you apply quotient rule ?

OpenStudy (nincompoop):

are you sure this is calculus 3?

OpenStudy (katielong):

i'm not sure, could you help? and yeah its Core 3 AQA a level maths

OpenStudy (nincompoop):

what is quotient?

OpenStudy (katielong):

the quotient rule?

OpenStudy (nincompoop):

well I asked if you can apply quotient rule and you said you don't know so I asked what is quotient

OpenStudy (katielong):

im really confused... so should i apply the quotient rule?

OpenStudy (katielong):

sorry im not making this easy for myself

OpenStudy (nincompoop):

\[\frac{p}{q} = pq^{-1}\] this is quotient, correct? do you see this format for you to apply quotient rule? no

OpenStudy (katielong):

the quotient rule: (g(x)f'(x) + g'(x)f(x) ) /g ^2

OpenStudy (nincompoop):

so that leaves us chain and product rule let us start with product \[a*b \] is product, so we can think a = (6x-3) and b = (2x-1)^5

OpenStudy (katielong):

okay

OpenStudy (nincompoop):

so what about chain rule?

OpenStudy (katielong):

da/dx =-3 (sorry its (6-3x) in the first bracket

OpenStudy (nincompoop):

it doesn't matter we can deal with specifics later, we are just trying to understand what rules we can possibly apply by looking at the formats

OpenStudy (katielong):

oh okay yeah

OpenStudy (katielong):

the chain rule: dy/dx=du/dx * dy/du

OpenStudy (katielong):

(which can be extended by adding more functions)

OpenStudy (nincompoop):

so we can basically apply both product and chain rule but we need to apply chain rule first

OpenStudy (katielong):

right.. soo..

OpenStudy (nincompoop):

which one would it be?

OpenStudy (nincompoop):

do you know power rule?

OpenStudy (katielong):

no we havent covered it but i watched a university lecture on this topic and it said d/dx (f(x))^n = nf^(n-1) df/dx

OpenStudy (katielong):

if that makes any sense?

OpenStudy (nincompoop):

yes

OpenStudy (katielong):

so the function = the differentiation * the integration or something?

OpenStudy (nincompoop):

integration? no

OpenStudy (katielong):

okay so what does the power rule state then?

OpenStudy (nincompoop):

this is why I am asking if you know power rule first, it is important that you know this very well first

OpenStudy (katielong):

oh okay

OpenStudy (nincompoop):

so do you?

OpenStudy (katielong):

nope

OpenStudy (nincompoop):

\[f(x) = x^{n} \rightarrow f'(x) = nx^{n-1}\]

OpenStudy (katielong):

is that the power rule? oh okay yeah i understand that one

OpenStudy (nincompoop):

okay we had x^n what if we had (2x+3)^n? this is where chain rule comes in first we try and do the power rule of everything in the parenthesis n(2x+3)^(n-1) then multiply it to the derivative of whatever is inside the parenthesis n(2x+3)^(n-1) * (2)

OpenStudy (katielong):

okay yeah im with you on that

OpenStudy (nincompoop):

so going back with your problem, you can apply the chain rule first before you can apply the product rule

OpenStudy (katielong):

so applying this to the question we get: 10(2x+3)^4

OpenStudy (nincompoop):

y=(6x-3)(2x-1)^5 chain rule with (2x-1)^5 then whatever answer you get, apply a product rule with 6x-3

OpenStudy (katielong):

=f(x)g'(x) + g(x)f'(x)

OpenStudy (nincompoop):

something along those lines

OpenStudy (katielong):

=(6x-3)(10(2x+3)^4) + (2x-1)^5 (-3) ?

OpenStudy (nincompoop):

what is the result of your chain rule?

OpenStudy (katielong):

* my mistake it should be (2x-1) not (2x-3) in the first set of brackets

OpenStudy (katielong):

would i expand the brackets?

OpenStudy (katielong):

=(60x-30)(2x-1)^4 + 32x *-x *-3 =

OpenStudy (nincompoop):

you lost me

OpenStudy (nincompoop):

what is the result of your chain rule?

OpenStudy (katielong):

I've lost myself.. aha i dont know?

OpenStudy (katielong):

10(2x-1)^4

OpenStudy (nincompoop):

why 10?

OpenStudy (nincompoop):

you're only bringing down 5

OpenStudy (nincompoop):

5(2x-1)^4 * (2)

OpenStudy (katielong):

oh yeah, but you multiply by d(2x-1)/dx which equals 2, therefore 2*5=10??

OpenStudy (nincompoop):

oh okay

OpenStudy (katielong):

is that correct?

OpenStudy (nincompoop):

yes

OpenStudy (katielong):

okay good, so now i presume we need to use the product rule?

OpenStudy (nincompoop):

yes

OpenStudy (katielong):

which would be (6-3x)*(the answer from before) + (2-1)^5 * d(6-3x)/dx

OpenStudy (katielong):

just that alone?

OpenStudy (nincompoop):

(6x-3) * 10(2x-1)^4

OpenStudy (nincompoop):

like that?

OpenStudy (katielong):

so... = 60x-30 * (2x-1)^4 ?

OpenStudy (katielong):

but that would only be f(x)g'(x) as oppposed to f(x)g'(x) + f'(x)g(x) ?

OpenStudy (nincompoop):

I think you're doing a lot of typos

OpenStudy (nincompoop):

yeah product rule is u'v+uv'

OpenStudy (katielong):

yeah.. thats an easier format!

OpenStudy (katielong):

uv' = (6x-3) * 10(2x-1)^4

OpenStudy (nincompoop):

go on

OpenStudy (katielong):

and u'v = ....

OpenStudy (katielong):

u'v = (2x-1)^5 * -3 ?

OpenStudy (nincompoop):

u = 6x-3 v = (2x-1)^5 u'v + uv' u'v = 6(2x-1)^5 uv' = (6x-3) 10(2x-1)^4

OpenStudy (nincompoop):

clear?

OpenStudy (katielong):

oh right yeah, sorry its (6-3x) in the brackets not (6x-3) but yes i understand the concept

OpenStudy (nincompoop):

laughing out loud okay

OpenStudy (nincompoop):

u = -3x+6 then?

OpenStudy (katielong):

.... therefore the grand answer is... 6(2x-1)^5 + (6x-3) 10(2x-1)^4 am i right?

OpenStudy (nincompoop):

sometimes, it might be less confusing if you did this list your u and v then list your u' and then v'

OpenStudy (katielong):

yeah, we did several questions and that was generally how i formatted them in my answers

OpenStudy (nincompoop):

then you follow all the formats of product, quotient and chain rules

OpenStudy (katielong):

aye i see

OpenStudy (nincompoop):

this kind of systematic approach will help you in calculus 2

OpenStudy (katielong):

just to complicate things.. the answer in the book says (2x-1)^4 (63-36x)

OpenStudy (nincompoop):

we are multiply buy a negative so there must be a negative somewhere

OpenStudy (katielong):

that's what i thought, there must be a negative

OpenStudy (katielong):

and a power of 5 due to it being uv'

OpenStudy (katielong):

but one last question, y=sin3xcos2x a simple product rule question right? but what is dsin3x/dx ?

OpenStudy (katielong):

is it 3 cos x ?

OpenStudy (nincompoop):

3 is constant and derivative of sin is cos

OpenStudy (katielong):

its been a while since i've done maths before today!

OpenStudy (katielong):

oh is it just cosx

OpenStudy (katielong):

oh thanks for that!

OpenStudy (nincompoop):

you're welcome I use product as u'v+uv' and quotient as (u'v - uv')/v^2 see the format? it's simple and no alternating you follow a simple alphabetical order